OpenStudy (aveline):

Prepare 100ml of an acetic acid and sodium acetate buffer w/ a pH of 4.50. HC2H3O2 (aq) = 0.53M NaC2H3O2 (s) = 136.08g/mol Ka HC2H3O2 = 1.8 * 10^-5

OpenStudy (aveline):

So far, this is what I've figured out: \[4.50=4.7+\log \frac{ A ^{-} }{ 0.53M }\] \[-0.20=\log \frac{ A ^{-} }{ 0.53M }\] \[0.33=A ^{-}\] So I need 0.33 moles of the sodium acetate and 0.53 of the acetate to make a solution with a pH of 4.50. Now I need to change these values so I can make 100ml of the buffer. How can I do that?

OpenStudy (aveline):

Does density have anything to do with this?

OpenStudy (aveline):

@AloneS @sweetburger

OpenStudy (aveline):

@ganeshie8

OpenStudy (aveline):

@Photon336

OpenStudy (aveline):

Anyone?